Night report 2021-08-25

Observers: Priscila Pessi

Support: Claudia Gutiérrez

TAT: Michael Fulton, Takashi Nagao, Jacob Teffs

EFOSC Calibrations:

SOFI Calibrations:

Weather:

Start: humidity ~%, wind ~m/s, seeing ~"

Middle: humidity ~%, wind ~m/s, seeing ~"

End: humidity ~%, wind ~m/s, seeing ~"

Comments (times in UT):

Telescope remained closed due to ice in the dome